• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 08:07:22 on November 3,2008

    Entry number 246407

    keyword=Owl Shift Summary


    Owl Shift Summary

    12:00 Shift started with no beam available.

    1:00 The MCC called to inform us that they are ready to send a tune beam to the hall.

    1:30 The MCC are still trying to set the requested beam position.

    1:45 Asked the MCC to do harp scans to make sure that the beam X and Y RMS are within the required limits (100-200um).

    2:40 The MCC Crew Chief called to let us know that the energy monitor is now working.

    2:50 The Q3 He level went below 60% suddenly which caused Q3 to trip.

    3:05 MCC finished the harp scan analysis. The X and Y RMS are within the required limits for 1H03A (1H03B is not working).

    3:35 ROCs 8, 9 and 10 stopped responding. Rebooted them and they are back up.

    3:40 Moved the target to the Reference Cell (H2 at 100 psgi).

    3:50 Could not do the rest of the BigBite wire chambers parameter check from last night (HV 1700 and 1750V). They trip too much! Xin will try this check later

    3:55 Moved the target to He3 Cell (Step #3 on the whiteboard).

    4:00 changed the raster size to 3x3, 4x4, 5x5, and 6x6mm to check the T3 rates.

    4:25 Moved the target to pickup coils position.

    4:30 Doing an NMR measurement. The polarization is 47%. The next NMR measurement is due in 2 hours (6:30AM).

    4:45 Moved the target back to the He3 Cell (Step #6 on the whiteboard).

    4:50 Check the T3 rates again at the 3x3 and 4x4 raters and found them to be the same (T3 = 8kHz).

    5:10 Started the production runs (finally!). We will take 2 half hour runs for Longitudinal

    5:45 Noticed that the DATAMON is not visible on the DAQ screen. Started it again and found that the deadtime is varying randomly between negative and positive large numbers (-20 to +60%). We stopped the bigbite cerenkov (check halog entry 246383) but the deadtime was still high.

    5:55 Q3 tripped again because of another helium level drop to 54%. Will wait until it is more than 60% to recycle Q3.

    6:15 Q3 is back up.

    6:20 Paged the run coordinator about the deadtime problem and we tried to pin the problem down on the phone. He will be here shortly to check this closely.


    7:00 Moved the target to the pickup coils position to take another NMR measurement The polarization is 42%.

    7:10 Moved the target to the Reference Cell (Nitrogen). We are going to take 7 runs at several gas pressure (step # 7 on the whiteboard).

    7:10 MCC having problems with beam.

    7:25 Beam is back.

    7:30 The run coordinator is here to check the high deadtime problem.

    7:45 The run coordinator said that we should stop running until we figure out the cause of the high dead time.

    8:00 Shift ends.